Shift report 20011030 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 30 2001 - 00:00:43 EST

  • Next message: Apache: "Shift report 20011030 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Oct 30 00:00:43 2001, Shift: 16:00-24:00
    Supervisor: CEJ, with: 
    Log book#10 pages 69 - 71
    
    Shift Summary:
    ------------------------------------------------------------
    1. Short test store with low lum 16:30-17:00. Several 
       attempts to ramp. Succes at 19:30. Nice store that
       lasted the rest of the shift. Approx 4.5 hours of 
       physics running.
    
    2. Runs: 
       5565:   7K tr-6 (Test store)
       5566:  97K tr-6 
       5567:  11K tr-6 (Magnets on)
       5569:  80K tr-6
       5572:  30K tr-6 (zero field run)
       5573:  ??K tr-6 (fs@8deg 1/5 B, still running)
      
    3. Magnet current measurement problem (meas=0) in the end 
       of run 5566 and in run 5567. MCR send a magnet expert 
       that fixed the problem. The magnets were on all the time!
       DC anodes tripped at the end of 5567.  
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 30 2001 - 00:01:41 EST